Assume that, on January 1, 2021, Sosa Enterprises paid $2,060,000 for its investment in 30,000 shares of Orioles Co. Further,Multiple Choice o $2.226,667. o $2210,000. o $2,197,250. o $2,074,250.
Answer #1

Solution:

Answer = $2,074,250

Percentage of holding held by Sosa enterprise in orioles.co. = 30,000 / 100,000 = 0.30 or 30%

The amount that would be reported by Sosa enterprise for the year ending 2021 balance sheet for its investment in Orioles.co. = Amount paid for investment + 30% of net income - 30% of dividends paid - 30% of additional depreciation.

= $2,060,000 + (30% \times $500,000) - (30% \times $300,000) - (30% \times $152,500)

= $2,060,000 + $150,000 - $90,000 - $45,750

= $2,074,250.

Explanation:

Depreciation assets = Total fair value - identifiable assets - amount belongs to land

Depreciation assets = $10,000,000 - $7,080,000 - $1,700,000

Depreciation assets = $1,220,000

Depreciation amount = $1,220,000 / 8 years = $152,500
Answer #2
The answer is $2,074,250.
answered by: anonymous